OpenStudy (anonymous):

hi, can you please help me. I am confused about a few problems. x(4x+24)=0

OpenStudy (azureilai):

you distribute the x so it becomes 4x^2+24x = 0

OpenStudy (anonymous):

the answer will have 2 option: x=0 and 4x+24=0 4(x+6)=0 x+6=0 x=-6 the answer are:x=0 x=-6